OpenStudy (anonymous):

(3x^2-y^2)dx+(xy-x^3y^{-1})dy=0 What technique is this??

OpenStudy (anonymous):

i tried separation, integrated factor, exact, bernelli

OpenStudy (anonymous):

is this implicit diferentiation

OpenStudy (anonymous):

i am just solving for y

OpenStudy (anonymous):

this is in linear algebra/diff equations

OpenStudy (anonymous):

i guess that this was implicit?

OpenStudy (anonymous):

is this a problem in a book?

OpenStudy (anonymous):

we have several techniques, Sepration D.E Integrated Factor, Homegeneous sub, F(ax+by) linear factor, bernalli, exact or second order D.E

OpenStudy (anonymous):

Its for homework, our professor made it

OpenStudy (anonymous):

siths must be typing out a real duesy

OpenStudy (anonymous):

i hope so ^.^

OpenStudy (anonymous):

Typically, given an ODE of this form, you want to check for exactness. \[\begin{align*} \underbrace{(3x^2-y^2)}_{M(x,y)}\,dx+\underbrace{\left(xy-\frac{x^3}{y}\right)}_{N(x,y)}\,dy&=0 \end{align*}\] which has partial derivatives \[\begin{cases} \dfrac{\partial M}{\partial y}=-2y\\\\ \dfrac{\partial N}{\partial x}=y-\dfrac{3x^2}{y} \end{cases}\] Since the partial derivatives aren't equal, the equation is not exact, but it's possible that an integrating factor exists, either a function of \(x\) or \(y\) or even a function of both. This doesn't look like a Bernoulli equation to me. Solving for the derivative gives \[\frac{dy}{dx}=\frac{y^2-3x}{xy-\dfrac{x^3}{y}}\] but you can't (as far as I can tell) write this in a Bernoulli form, i.e. \[\frac{dy}{dx}+f(x)y+g(x)y^n=0\]

OpenStudy (anonymous):

@mattyboyy

OpenStudy (anonymous):

We might as well check to see if the function is homogeneous (I doubt it, but it doesn't hurt to check). An ODE \(\dfrac{dy}{dx}=f(x,y)\) is homogeneous if the function \(f\) satisfies \[f(tx,ty)=t^kf(x,y)\] for some real \(k\). \[\begin{align*} f(tx,ty)&=\frac{t^2y^2-3tx}{(tx)(ty)-\dfrac{t^3x^3}{ty}}\\\\ &=\frac{t^2y^2-3tx}{t^2xy-t^2\dfrac{x^3}{y}}\\\\ &=t\frac{ty^2-3x}{txy-\dfrac{x^3}{y}}\\\\ &\neq t^kf(x,y) \end{align*}\] and so the equation is indeed not homogeneous.

OpenStudy (anonymous):

So do you think we need to try the not exact? put it in the form My-Mx/N to make a new equation and then find the integrating factor?

OpenStudy (anonymous):

siths going hard with the equation button lol

OpenStudy (anonymous):

Yeah and it is awesome!

OpenStudy (anonymous):

We have two options for a single-variable integrating factor: \[\mu(x)=\exp\left(\int\frac{M_y-N_x}{N}\,dx\right)\quad\text{or}\quad\mu(y)=\exp\left(\int\frac{N_x-M_y}{M}\,dy\right)\] It looks like the first one will suffice. \[\begin{align*}\mu(x)&=\exp\left(\int\frac{-2y-\left(y-\dfrac{3x^2}{y}\right)}{xy-\dfrac{x^3}{y}}\,dx\right)\\\\ &=\exp\left(-3\int\frac{y-\dfrac{x^2}{y}}{xy-\dfrac{x^3}{y}}\,dx\right)\\\\ &=\exp\left(-3\int\frac{y-\dfrac{x^2}{y}}{x\left(y-\dfrac{x^2}{y}\right)}\,dx\right)\\\\ &=\exp\left(-3\int\frac{dx}{x}\right)\\\\ &=\exp\left(-3\ln|x|\right)\\\\ &=\exp\left(\ln|x|^{-3}\right)\\\\ \mu(x)&=\frac{1}{x^3} \end{align*}\]

OpenStudy (anonymous):

wow your algebra skills are awesome

OpenStudy (anonymous):

Thank you so much!!

OpenStudy (anonymous):

@SithsAndGiggles College?

OpenStudy (anonymous):

Yes. Also, I'm not using the Equation editor. All manual. @edcion You're welcome!

OpenStudy (anonymous):

It would have been a lot more work if I'd used the Eq button :P

OpenStudy (anonymous):

Very interesting, you must have a lot of practice

OpenStudy (anonymous):

after multiplying by the integrated factor the two terms are now Exact so it works thank you again!

OpenStudy (anonymous):

You're welcome!